# Optimality of quasi-Monte Carlo methods and suboptimality of the sparse-grid Gauss–Hermite rule

### v1

@@ 1 Introduction
$I(f):=\int_{\mathbb{R}^{d}}f(\boldsymbol{x})\frac{\mathrm{e}^{-|\boldsymbol{x}|^{2}/2}}{(2\pi)^{d/2}}\mathrm{d}\boldsymbol{x}\approx Q_{N}(f),$
@@
where $|\boldsymbol{x}|=\sqrt{x_{1}^{2}+\dotsb+x_{d}^{2}}$ for $\boldsymbol{x}\in\mathbb{R}^{d}$ is the Euclidean norm, and $Q_{N}$ is a suitable numerical integration rule using $N$ evaluations of $f$. Such problems arise across a broad range of applications, including data assimilation [36], robotics [3], life insurance [23], physics [45]. They also serve as simpler problems to which more complex integration tasks can be reduced [20].
@@
In high dimensions $d\gg 1$, two popular classes of numerical integration rules $Q_{N}$ are sparse-grid quadrature [6] and quasi-Monte Carlo (QMC) methods [13]. Sparse-grid quadrature is a numerical integration technique based on univariate quadrature rules. Unlike fully tensorized quadrature, it significantly reduces the number of evaluation points, making it attractive for high-dimensional problems. A commonly used univariate quadrature for approximating $I(f)$ is the Gauss–Hermite rule [38, 56, 19, 4, 66], which is the focus of this paper. Quasi-Monte Carlo, on the other hand, is an equal-weight numerical integration. Its core idea is to carefully choose the quadrature points to improve accuracy. Examples of point sets include low-discrepancy point sets and sequences, lattice rules, and digital nets; these classes are not mutually exclusive. These methods are also widely used for integration with respect to the Gaussian measure and have found applications across a diverse range of scientific fields, such as uncertainty quantification [9, 42], computational physics [37], finance [24, 43], and machine learning [1, 47].
@@
In this paper, we show that sparse-grid Gauss–Hermite quadrature is suboptimal, whereas several quasi-Monte Carlo methods are optimal in terms of convergence rate. More precisely, for $L^{2}$-Sobolev spaces of degree $\alpha$, the sparse-grid Gauss–Hermite rule achieves the convergence rate $O(N^{-\alpha/2})$; several quasi-Monte Carlo (QMC) methods together with change of variables achieve $O(N^{-\alpha})$ up to a logarithmic factor. The rate $O(N^{-\alpha})$ is in fact optimal, as there exists a matching lower bound established by Dick et al. [11]. In this sense, QMC methods considered herein are rate-optimal. In contrast, we also show that the rate $O(N^{-\alpha/2})$ for the sparse-grid Gauss–Hermite rule likewise unimprovable, by proving a corresponding lower bound.
@@
Such a gap in convergence rates between the sparse-grid Gauss–Hermite rule and QMC methods has also been observed numerically by Dick et al. [11] and in subsequent work by Nuyens and one of the present authors [55]. Our theoretical results provide a rigorous explanation of these empirical findings.
@@
Analysis of the sparse-grid Gauss–Hermite rule and QMC methods on the whole space has been carried out before, with many papers set in the context of partial differential equations with random coefficients [7, 15, 17, 21, 30, 31, 32, 40, 49, 58]. This paper provides theoretical results for a standard function space, without specific structure of the underlying model problem such as the form of the random field or partial differential equations whose differential operator is given by such a field. In the same vein, aforementioned work by Dick et al. [11] studied quasi-Monte Carlo methods in a Sobolev space, the space we also consider herein. They showed that the QMC method considered there attains the optimal convergence rate by proving matching upper and lower bounds, although the logarithmic factor was not optimal. Nuyens and one of the present authors considered a rank-$1$ lattice QMC over $\mathbb{R}^{d}$ [55] for functions with suitable decay. We revisit their results and show that the resulting QMC method achieves the optimal rate in the Sobolev space considered herein, although the logarithmic factor is again not optimal. In [64], Hyvönen, Karvonen, and one of the present authors considered a change of variables that maps functions on the real line to the unit interval. Although such an approach typically incurs undesirable unboundedness near the boundary, they showed that, with a carefully chosen Möbius transformation, the resulting mapping sends univariate functions from the Gaussian Sobolev space on the real line to the periodic Sobolev space of the same smoothness on the unit interval. In this paper, by applying such a transformation component-wise to $d$-variate functions, we show that the resulting function belongs to two classes of Sobolev spaces that are important in QMC: Korobov spaces and unanchored Sobolev spaces. Using this, we establish that QMC rules with the Möbius transformation inherit the convergence rates of the respective spaces. These rates turn out to be optimal in Gaussian Sobolev spaces: rank-$1$ lattice rules achieve the optimal rate up to a logarithmic factor, whereas higher-order digital nets achieve the optimal rate including the logarithmic factor.
@@
Dũng and Nguyen [16] proposed two numerical integration methods that are optimal – including the logarithmic factor – in $L^{p}$-Sobolev spaces for $1<p<\infty$. One method requires implementing a partition of unity with suitable properties. The other method appears to be implementable, but the quadrature weights depend on various indices. The results in this paper show that equal-weight quadrature rules suffice to obtain an optimal numerical integration rule.
@@
The remainder of the paper is organized as follows. In Section 2 we introduce necessary concepts and state preliminary results. Sections 3 and 4 present our main results on the sparse-grid Gauss–Hermite quadrature and QMC methods, respectively. These results build on the authors’ previous works [29, 41, 64]. Finally, Section 5 concludes the paper.
@@ 2.1 The Sobolev space
where $\rho(\boldsymbol{x})=\frac{\mathrm{e}^{-|\boldsymbol{x}|^{2}/2}}{(2\pi)^{d/2}}$ and $D^{\mathbf{r}}f$ is the weak derivative of $f$. Here, $\|\cdot\|_{L_{\rho}^{2}(\mathbb{R}^{d})}$ is the norm induced by the weighted $L^{2}$-inner product $\langle f,g\rangle_{L_{\rho}^{2}}:=\int_{\mathbb{R}^{d}}f(\boldsymbol{x})g(\boldsymbol{x})\rho(\boldsymbol{x})\mathrm{d}\boldsymbol{x}$ for $f,g\in L_{\rho}^{2}(\mathbb{R}^{d})$.
It is known that this space and the so-called Hermite space coincide as vector spaces, with equivalent norms. More specifically, Dick et al. [11] studied quasi-Monte Carlo methods in the Hermite space and showed that every element of this space also belongs to $H_{\rho}^{\alpha}(\mathbb{R}^{d})$, with the embedding being continuous; the reverse inclusion is discussed in [41, 16]. This space is also known to coincide with the so-called modulation space, as discussed in [18]. Further properties of Hermite spaces are discussed in [25, 26].
@@
Hence, any algorithm achieving this convergence rate $\frac{(\ln N)^{\frac{d-1}{2}}}{N^{\alpha}}$ may be regarded as optimal, in terms of the worst-case error convergence rate. We will show that several QMC methods attain a rate of the form $\frac{(\ln N)^{\gamma}}{N^{\alpha}}$, where the exponent $\gamma$ depends on the method. The rate $\gamma=\frac{d-1}{2}$ turns out to be attainable, hence optimal in the sense above. In contrast, the sparse-grid quadrature based on the Gauss–Hermite rule is of the rate $N^{-\alpha/2}$, up to a logarithmic factor, which is subotimal in the sense of worst-case error convergence rate.
Note that [11, Theorem 1] is formulated for linear algorithms. As noted in [11, Section 3], however, in view of for example [2] (see also [53, Section 4.2]), their lower bound also holds for nonlinear algorithms. Moreover, the implied constant in the result above may differ from that in [11, Theorem 1], since the present bound is formulated for the Sobolev space, whereas [11, Theorem 1] considers a norm-equivalent space, namely the Hermite space.
@@
Throughout the paper, for mappings $f,g\colon S\to[0,\infty]$ on any set $S$, we write $f\lesssim g$ if there exists a constant $c\in(0,\infty)$ such that $f(x)\leq cg(x)$ holds for all $x\in S$. We write $f\asymp g$ if both $f\lesssim g$ and $g\lesssim f$. If the constant may depend on specific parameters, we indicate this by subscripts. For example, $f\lesssim_{d}g$ means there exists a constant $c_{d}\in(0,\infty)$, which may depend on $d$, such that $f(x)\leq c_{d}\,g(x)$ for all $x\in S$; likewise, $f\asymp_{d}g$ means $f\lesssim_{d}g$ and $g\lesssim_{d}f$.
@@ 2.2 Tensorised operators
Proof.
Fix $f\in H_{\rho}^{\alpha}$. Following the argument in [41, Lemma 2.1] we see
$\langle D^{\mathbf{r}}f,H_{\mathbf{k}}\rangle_{L_{\rho}^{2}}=\biggl(\prod_{j=1}^{d}\prod_{m=1}^{\mathrm{r}_{j}}\bigl(k_{j}+m\bigr)\biggr)^{1/2}\langle f,H_{\mathbf{k}+\mathbf{r}}\rangle_{L_{\rho}^{2}}\quad\text{for }\text{$\mathbf{k}$}\in\mathbb{N}_{0}^{d},\ \mathbf{r}\in\{0,\dots,\alpha\}^{d}.$
@@
$\left[T_{1}\otimes\dotsb\otimes T_{d}\Bigg(\prod_{j=1}^{d}H_{k_{j}}\Bigg)\right](\boldsymbol{x})=\prod_{j=1}^{d}[T_{j}(H_{k_{1}})](x_{j})\qquad\text{for }\prod_{j=1}^{d}H_{k_{j}}\in\mathcal{P}$
and extending it linearly to $\mathcal{P}$.
Then, using the argument for tensor product Hilbert spaces [57, Section VIII.10] we see $\sup_{p\in\mathcal{P},\|p\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}=1}\|T_{1}\otimes\dotsb\otimes T_{d}(p)\|=\prod_{j=1}^{d}\|T_{j}\|_{H_{\rho}^{\alpha}(\mathbb{R})\to H_{\rho}^{\alpha}(\mathbb{R})}$ on $\mathcal{P}$, and from Proposition 2.2, the operator $T_{1}\otimes\dotsb\otimes T_{d}$ extends continuously to $H_{\rho}^{\alpha}(\mathbb{R}^{d})$ with the same operator norm.
Note that any continuous linear functional on $H_{\rho}^{\alpha}(\mathbb{R}^{d})$ is a bounded linear operator in our setting. Indeed, since the total measure is normalised to $1$, any constant belongs to $H_{\rho}^{\alpha}(\mathbb{R}^{d})$, and if $T_{j}$ is a functional then we have $T_{j}(f)\in H_{\rho}^{\alpha}(\mathbb{R})$ with $\|T_{j}(f)\|_{H_{\rho}^{\alpha}(\mathbb{R})}=|T_{j}(f)|$ for $f\in H_{\rho}^{\alpha}(\mathbb{R})$. Operators with the symbol $\otimes$ in this paper are constructed explicitly in this manner.
@@ 2.3 Sobolev embedding for spaces with dominating mixed smoothness
$\sup_{\boldsymbol{x}\in\mathbb{R}^{d}}|v(\boldsymbol{x})|$ $\leq\|v\|_{W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})},$ (1)
@@
$|v(\boldsymbol{x})-v(\boldsymbol{y})|$ $\leq d^{3/4}\|v\|_{W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})}|\boldsymbol{x}-\boldsymbol{y}|_{2}^{1/2}\qquad\text{for all }\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}.$ (2)
Proof.
We will first prove the statement for infinitely differentiable functions with compact support, $C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$, and then extend it by a density argument. For $v\in C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$, following [5, Proof of Theorem 8.8] we get
@@
which shows the second inequality for $C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$. Using the density of $C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$ completes the proof. ∎

### v2

@@ Abstract
keywords:
high-dimensional integration, quasi-Monte Carlo, digital net, rank-1 lattice rule, Möbius transformation, Smolyak algorithm, Gauss–Hermite quadrature, optimal algorithm
{MSCcodes}
65C05, 65D30, 65D32, 65D40
@@ 1 Introduction
$I(f):=\int_{\mathbb{R}^{d}}f(\boldsymbol{x})\frac{\mathrm{e}^{-|\boldsymbol{x}|_{{2}}^{2}/2}}{(2\pi)^{d/2}}\mathrm{d}\boldsymbol{x}\approx Q_{N}(f),$
@@
where $|\boldsymbol{x}|_{2}=\sqrt{x_{1}^{2}+\dotsb+x_{d}^{2}}$ for $\boldsymbol{x}\in\mathbb{R}^{d}$ is the Euclidean norm, and $Q_{N}$ is a suitable numerical integration rule using $N$ evaluations of $f$. Such problems arise across a broad range of applications, including data assimilation [undefak], robotics [undefb], life insurance [undefw], physics [undefau]. They also serve as simpler problems to which more complex integration tasks can be reduced [undeft].
@@
In high dimensions $d\gg 1$, two popular classes of numerical integration rules $Q_{N}$ are sparse-grid quadrature [undefe] and quasi-Monte Carlo (QMC) methods [undefm]. Sparse-grid quadrature is a numerical integration technique based on univariate quadrature rules. Unlike fully tensorized quadrature, it significantly reduces the number of evaluation points, making it attractive for high-dimensional problems. A commonly used univariate quadrature for approximating $I(f)$ is the Gauss–Hermite rule [undefam, undefaaf, undefs, undefc, undefaap], which is the focus of this paper. Quasi-Monte Carlo, on the other hand, is an equal-weight numerical integration. Its core idea is to carefully choose the quadrature points to improve accuracy. Examples of point sets include low-discrepancy point sets and sequences, lattice rules, and digital nets; these classes are not mutually exclusive. These methods are also widely used for integration with respect to the Gaussian measure and have found applications across a diverse range of scientific fields, such as uncertainty quantification [undefi, undefaq], computational physics [undefal], finance [undefx, undefar], and machine learning [undef, undefaw].
@@
In this paper, we show that sparse-grid Gauss–Hermite quadrature is suboptimal, whereas several quasi-Monte Carlo methods are optimal in terms of convergence rate. More precisely, for $L^{2}$-Sobolev spaces of degree $\alpha$, the sparse-grid Gauss–Hermite rule achieves the convergence rate $O(N^{-\alpha/2})$; several quasi-Monte Carlo (QMC) methods together with change of variables achieve $O(N^{-\alpha})$ up to a logarithmic factor. The rate $O(N^{-\alpha})$ is in fact optimal, as there exists a matching lower bound established by Dick et al. [undefk]. In this sense, QMC methods considered herein are rate-optimal. In contrast, we also show that the rate $O(N^{-\alpha/2})$ for the sparse-grid Gauss–Hermite rule likewise unimprovable, by proving a corresponding lower bound.
@@
Such a gap in convergence rates between the sparse-grid Gauss–Hermite rule and QMC methods has also been observed numerically by Dick et al. [undefk] and in subsequent work by Nuyens and one of the present authors [undefaae]. Our theoretical results provide a rigorous explanation of these empirical findings.
@@
Analysis of the sparse-grid Gauss–Hermite rule and QMC methods on the whole space has been carried out before, with many papers set in the context of partial differential equations with random coefficients [undefg, undefo, undefq, undefu, undefad, undefae, undefaf, undefao, undefay, undefaah]. This paper provides theoretical results for a standard function space, without specific structure of the underlying model problem such as the form of the random field or partial differential equations whose differential operator is given by such a field. In the same vein, aforementioned work by Dick et al. [undefk] studied quasi-Monte Carlo methods in a Sobolev space, the space we also consider herein. They showed that the QMC method considered there attains the optimal convergence rate by proving matching upper and lower bounds, although the logarithmic factor was not optimal. Nuyens and one of the present authors considered a rank-$1$ lattice QMC over $\mathbb{R}^{d}$ [undefaae] for functions with suitable decay. We revisit their results and show that the resulting QMC method achieves the optimal rate in the Sobolev space considered herein, although the logarithmic factor is again not optimal. In [undefaan], Hyvönen, Karvonen, and one of the present authors considered a change of variables that maps functions on the real line to the unit interval. Although such an approach typically incurs undesirable unboundedness near the boundary, they showed that, with a carefully chosen Möbius transformation, the resulting mapping sends univariate functions from the Gaussian Sobolev space on the real line to the periodic Sobolev space of the same smoothness on the unit interval. In this paper, by applying such a transformation component-wise to $d$-variate functions, we show that the resulting function belongs to two classes of Sobolev spaces that are important in QMC: Korobov spaces and unanchored Sobolev spaces. Using this, we establish that QMC rules with the Möbius transformation inherit the convergence rates of the respective spaces. These rates turn out to be optimal in Gaussian Sobolev spaces: rank-$1$ lattice rules achieve the optimal rate up to a logarithmic factor, whereas higher-order digital nets achieve the optimal rate including the logarithmic factor.
@@
Dũng and Nguyen [undefp] proposed two numerical integration methods that are optimal – including the logarithmic factor – in $L^{p}$-Sobolev spaces for $1<p<\infty$. One method requires implementing a partition of unity with suitable properties. The other method appears to be implementable, but the quadrature weights depend on various indices. The results in this paper show that equal-weight quadrature rules suffice to obtain an optimal numerical integration rule.
@@
The remainder of the paper is organized as follows. In Section 2 we introduce necessary concepts and state preliminary results. Sections 3 and 4 present our main results on the sparse-grid Gauss–Hermite quadrature and QMC methods, respectively. These results build on the authors’ previous works [undefac, undefap, undefaan]. Finally, Section 5 concludes the paper.
@@ 2.1 The Sobolev space
where $\rho(\boldsymbol{x})=\frac{\mathrm{e}^{-|\boldsymbol{x}|_{{2}}^{2}/2}}{(2\pi)^{d/2}}$ and $D^{\mathbf{r}}f$ is the weak derivative of $f$. Here, $\|\cdot\|_{L_{\rho}^{2}(\mathbb{R}^{d})}$ is the norm induced by the weighted $L^{2}$-inner product $\langle f,g\rangle_{L_{\rho}^{2}}:=\int_{\mathbb{R}^{d}}f(\boldsymbol{x})g(\boldsymbol{x})\rho(\boldsymbol{x})\mathrm{d}\boldsymbol{x}$ for $f,g\in L_{\rho}^{2}(\mathbb{R}^{d})$. Throughout this paper, we denote by $L_{\mathrm{loc}}^{1}(\mathbb{R}^{d})$ the space of locally integrable functions on $\mathbb{R}^{d}$.
It is known that this space and the so-called Hermite space introduced in [undefaj] and generalized in [undefk] coincide as vector spaces, with equivalent norms. More specifically, Dick et al. [undefk] studied quasi-Monte Carlo methods in the Hermite space and showed that every element of this space also belongs to $H_{\rho}^{\alpha}(\mathbb{R}^{d})$, with the embedding being continuous; the reverse inclusion is discussed in [undefap, undefp]. This space is also known to coincide with the so-called modulation space, as discussed in [undefr]. Further properties of Hermite spaces are discussed in [undefy, undefas, undefz].
@@
Hence, any algorithm achieving this convergence rate $\frac{(\ln N)^{\frac{d-1}{2}}}{N^{\alpha}}$ may be regarded as optimal, in terms of the worst-case error convergence rate. We will show that several QMC methods attain a rate of the form $\frac{(\ln N)^{\gamma}}{N^{\alpha}}$, where the exponent $\gamma$ depends on the method. The exponent $\gamma=\frac{d-1}{2}$ turns out to be attainable, hence optimal in the sense above. In contrast, the sparse-grid quadrature based on the Gauss–Hermite rule is of the rate $N^{-\alpha/2}$, up to a logarithmic factor, which is subotimal in the sense of worst-case error convergence rate.
Note that [undefk, Theorem 1] is formulated for linear algorithms. As noted in [undefk, Section 3], however, in view of for example [undefa] (see also [undefaac, Section 4.2]), their lower bound also holds for nonlinear algorithms. Moreover, the implied constant in the result above may differ from that in [undefk, Theorem 1], since the present bound is formulated for the Sobolev space, whereas [undefk, Theorem 1] considers a norm-equivalent space, namely the Hermite space. Further characterizations of Hermite spaces are investigated in [undefas].
@@
Throughout the paper, for mappings $f,g\colon S\to[0,\infty]$ on any set $S$, we write $f\lesssim g$ if there exists a constant $c\in(0,\infty)$ such that $f(x)\leq cg(x)$ holds for all $x\in S$. We write $f\asymp g$ if both $f\lesssim g$ and $g\lesssim f$. If the constant may depend on specific parameters, we indicate this by subscripts. For example, $f\lesssim_{d}g$ means there exists a constant $c_{d}\in(0,\infty)$, which may depend on $d$, such that $f(x)\leq c_{d}\,g(x)$ for all $x\in S$; likewise, $f\asymp_{d}g$ means $f\lesssim_{d}g$ and $g\lesssim_{d}f$. We also use the following notations: $\mathbf{a}!:=\prod_{j=1}^{d}a_{j}!$ and $\binom{\mathbf{a}}{\mathbf{b}}:=\prod_{j=1}^{d}\binom{a_{j}}{b_{j}}$ for $\mathbf{a},\mathbf{b}\in\mathbb{N}^{d}_{0}$. Furthermore, for $\mathbf{k}\in\mathbb{N}_{0}^{d}$ we let $|\mathbf{k}|:=k_{1}+\dotsb+k_{d}$.
@@ 2.2 Tensorised operators
Proof 2.3.
Fix $f\in H_{\rho}^{\alpha}$. Following the argument in [undefap, Lemma 2.1] we see
$\langle D^{\mathbf{r}}f,H_{\mathbf{k}}\rangle_{L_{\rho}^{2}}=\biggl(\prod_{j=1}^{d}\prod_{m=1}^{\mathrm{r}_{j}}\bigl(k_{j}+m\bigr)\biggr)^{1/2}\langle f,H_{\mathbf{k}+\mathbf{r}}\rangle_{L_{\rho}^{2}}\quad\text{for }\mathbf{k}\in\mathbb{N}_{0}^{d},\ \mathbf{r}\in\{0,\dots,\alpha\}^{d}.$
@@
$\left[T_{1}\otimes\dotsb\otimes T_{d}\Bigg(\prod_{j=1}^{d}H_{k_{j}}\Bigg)\right](\boldsymbol{x})=\prod_{j=1}^{d}[T_{j}(H_{k_{{j}}})](x_{j})\qquad\text{for }\prod_{j=1}^{d}H_{k_{j}}\in\mathcal{P}$
and extending it linearly to $\mathcal{P}$. Then, using the argument for tensor product Hilbert spaces [undefaag, Section VIII.10] we see $\sup_{p\in\mathcal{P},\|p\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}=1}\|T_{1}\otimes\dotsb\otimes T_{d}(p)\|_{{H_{\rho}^{\alpha}(\mathbb{R}^{d})}}=\prod_{j=1}^{d}\|T_{j}\|_{H_{\rho}^{\alpha}(\mathbb{R})\to H_{\rho}^{\alpha}(\mathbb{R})}$ on $\mathcal{P}$, and from Proposition 2.2, the operator $T_{1}\otimes\dotsb\otimes T_{d}$ extends continuously to $H_{\rho}^{\alpha}(\mathbb{R}^{d})$ with the same operator norm.
Note that this construction applies also when the $T_{j}$ are bounded linear functionals. In such a case, since $T_{j}(f)$ is a constant function for $f\in H_{\rho}^{\alpha}(\mathbb{R})$ and the total measure is normalised to $1$, we have $\|T_{j}(f)\|_{H_{\rho}^{\alpha}(\mathbb{R})}=|T_{j}(f)|$. From this, we see that $T_{j}$ is also a bounded operator from $H_{\rho}^{\alpha}(\mathbb{R})$ to itself with $\|T_{j}\|_{H_{\rho}^{\alpha}(\mathbb{R})\to H_{\rho}^{\alpha}(\mathbb{R})}=\|T_{j}\|_{H_{\rho}^{\alpha}(\mathbb{R})\to\mathbb{R}}$. Similarly, for the functional $T_{1}\otimes\dotsb\otimes T_{d}$ on $\mathcal{P}$ constructed as above, $T_{1}\otimes\dotsb\otimes T_{d}(p)$ is a constant function so we have $|T_{1}\otimes\dotsb\otimes T_{d}(p)|=\|T_{1}\otimes\dotsb\otimes T_{d}(p)\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$. Hence, $T_{1}\otimes\dotsb\otimes T_{d}$ satisfies $\sup_{p\in\mathcal{P},\|p\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}=1}|T_{1}\otimes\dotsb\otimes T_{d}(p)|=\prod_{j=1}^{d}\|T_{j}\|_{H_{\rho}^{\alpha}(\mathbb{R})\to\mathbb{R}}$ and admits a continuous extension to $H_{\rho}^{\alpha}(\mathbb{R}^{d})$ with the same functional norm. Operators with the symbol $\otimes$ in this paper are constructed explicitly in this manner.
@@ 2.3 Sobolev embedding for spaces with dominating mixed smoothness
(1) $\sup_{\boldsymbol{x}\in\mathbb{R}^{d}}|v(\boldsymbol{x})|$ $\leq\|v\|_{W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})},$
@@
(2) $|v(\boldsymbol{x})-v(\boldsymbol{y})|$ $\leq d^{3/4}\|v\|_{W_{\mathrm{mix}}^{1,2}(\mathbb{R}^{d})}|\boldsymbol{x}-\boldsymbol{y}|_{2}^{1/2}\qquad\text{for all }\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}.$
Proof 2.5.
We will first prove the statement for infinitely differentiable functions with compact support, $C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$, and then extend it by a density argument. For $v\in C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$, following [undefd, Proof of Theorem 8.8] we get
@@
which shows the second inequality for $C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$. Using the density of $C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$ completes the proof.